This question is based on the following cash flows

C0=$10,000 I1=$2,000 I2=$2,000 I3=$3,000 I4=$3,000 I5=$2,500 L=$1,000
0 1 2 3 4 5

C: Cost, I: Income, L: Salvage

The escalation rate is 10% per year in this example.

Calculate the NPV of this investment assuming an escalated dollar minimum rate of return of 8%
Answer #1

Present value of investment = $10,000

Present value of income = 2,000 / (1+0.08) + 2,000 / (1+0.08)2 + 3,000 / (1+0.08)3 + 3,000 / (1+0.08)4 + 2,500 / (1+0.08)5

= $9,854.57

Present value of Salvage = 1,000 / (1+0.1)5

= $620.92

Net present value of investment = 9,854.57 + 620.92 - 10,000

= $475,49